A grocery store bought ice cream for $2.60 a half gallon and stored it in two freezers. During the night, one freezer malfunctioned and ruined 15 half gallons. If the remaining ice cream is sold for $3.98 per half gallon, how many half gallons did the store buy if it made a profit of $64.20?
Found 2 solutions by benni1013, lwsshak3:
Answer by benni1013(206)

A great problem dealing with profit margin. We know they purchased the ice cream at $2.60 and resold them at $3.98 creating a profit margin of $1.38/unit. To find the total that they sold during the day divide the profit margin divided by the profit made, $64.20/$1.38 or 46.52 or round up for 47. Add the addition 15 that were ruined to get 61 or 62 half gallons of ice cream.

Answer by lwsshak3(11628)

A grocery store bought ice cream for $2.60 a half gallon and stored it in two freezers. During the night, one freezer malfunctioned and ruined 15 half gallons. If the remaining ice cream is sold for $3.98 per half gallon, how many half gallons did the store buy if it made a profit of $64.20?
**
let x=number of half gallons of ice cream the store bought
x-15=number of unspoiled half gallons of ice cream
..
3.98(x-15)-2.60x=64.20
3.98x-59.70)-2.60x=64.20
1.38x=64.20+59.70=123.90
x=123.90/1.38≈89.78≈90 half gallons
number of half gallons of ice cream the store bought=90 half gallons